OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (anonymous):

distribute again the -3 on both sides

OpenStudy (anonymous):

idk what that means:/

OpenStudy (anonymous):

it means you multiply the number on the outside of the parentheses by all the numbers on the inside of them

OpenStudy (anonymous):

can you do it so i can see?

OpenStudy (anonymous):

-3x-9=-3x-3-5 then you need to simplify where you can

OpenStudy (anonymous):

idk how to simplify it

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

you end up with -3x-9 = -3x -8 then add 9 to each side and you get -3x = -3x + 1 these aren't equal so it is unsolvable

OpenStudy (nurali):

−3(x + 3) = −3(x + 1) − 5 -3x-9=-3x-3-5 -3x-9=-3x-8 -3x+3x=-8+9 0=1 No solution.

OpenStudy (anonymous):

@@

OpenStudy (anonymous):

Here is an example: -4(x + 4) = −2(x + 1) − 3 -4x + 4 = -2x - 2 - 3 +2x +2x -2x + 4 = -2-3 <-- Solve here -2x + 4 = -5 - 4 -2x = -9 /-2 /-2 x = 4.5